# Flipping the Dialogue: Training and Evaluating User Language Models

### v1

@@ Abstract
Conversations with LMs involve two participants: a human user leading the conversation, and an LM assistant responding to the user’s request. To satisfy this specific role, LMs are post-trained to be helpful assistants – optimized to produce exhaustive and well-structured responses, free of ambiguity and grammar errors. User utterances, on the other hand, are rarely perfected, with each user phrasing requests in unique ways, sometimes putting in partial effort at each turn and refining on the fly. To evaluate LM performance in realistic settings, prior work simulated users in multi-turn conversations, often prompting an LLM originally trained to be a helpful assistant to act as a user. However, we show that assistant LMs make for poor user simulators, with the surprising finding that better assistants yield worse simulators. Instead, we introduce purpose-built User Language Models (User LMs) - models post-trained to simulate human users in multi-turn conversations. Through various evaluations, we show how User LMs align better with human behavior and achieve better simulation robustness than existing simulation methods. When leveraging User LMs to simulate coding and math conversations, the performance of a strong assistant (GPT-4o) drops from 74.6% to 57.4%, confirming that more realistic simulation environments lead to assistant struggles as they fail to cope with the nuances of users in multi-turn setups.
@@ 2.1 Problem Description
Similar to how assistant LMs must follow instructions, user LMs must follow an intent that directs the conversation. Initial experimentation revealed that there exists a fine balance in defining user intent. On one extreme, not providing any intent at all reduces the usability of the user simulator, as it cannot be effectively directed towards a task. On the other extreme, a fully-specified intent that contains all information of the user utterances renders the simulator obsolete: it simply parrots information from the intent with little added value. In this work, we define user intents as high-level conversation objectives: capturing the overall goal of the user without mentioning explicit details (see examples in Figure 2) – achieving a balance between the two extremes. Preliminary analyses showed that user LMs trained on high-level intents were practically more useful than ones trained without intent or with fully-specified intents: they can be guided to simulate specific conversations, but are in charge of language choices that allow for diverse and realistic simulation.
@@ 2.2 Experimental Details
We use WildChat Zhao et al. (2024), which contains 478,498 English conversations between ChatGPT and users from 192 countries. We also noticed many near-duplicate samples and performed de-duplication, which resulted in 384,336 conversations (see details in Appendix A). The authors of WildChat did not define data splits (i.e., train vs. test); therefore, we created our own by first identifying unique users based on their hashed IP addresses and countries. For each country, we randomly split users into a 90/5/5 split for training, validation, and testing. All conversations from the same user were then assigned to the same split, ensuring that each split contains conversations from entirely distinct users. This resulted in 343,951 conversations for training, 22,442 for validation, and 17,943 for testing. After formatting conversations as we described in §2.1, we obtained 1,047,930 training samples, 118,291 validation samples, and 137,224 testing samples.
@@
We perform full-parameter fine-tuning of Llama3-8b-Base and Llama3.2-1b-Base. As we will show, we find that starting from a base model leads to better performance than an instruction-tuned model. We used a maximum sequence length of 2048 tokens, a batch size of 1024 samples, and a learning rate of 2e-5. Training was performed on four NVIDIA RTX A6000 GPUs, taking 62 hours to train the 1b model and 227 hours to train the 8b model. We refer to our fine-tuned user LMs as UserLM-1b and UserLM-8b.
@@
We compare to multiple assistant LMs: Llama3.2-1b-Instruct, Llama3-8b-Instruct, and GPT-4o. We prompted these assistant LMs to play the role of a human user and respond to the assistant given the conversation history so far, basing our prompt on recent studies Chang et al. (2025); Ivey et al. (2024). We gave explicit instructions on how to replicate user behavior: splitting content across turns, not being verbose, or making occasional typos, without overdoing such phenomena (see prompt in Appendix D). We also compare to USP-8b Wang et al. (2025), a fine-tuned version of Llama3-8b-Base on a subset of 94,874 LMSys-Chat conversations to generate user utterances Zheng et al. (2023). All baselines were instructed to respond with <|endconversation|> if they believe the user intent is fulfilled and that the conversation should end at this turn.
@@ 2.3 Initial Analyses
We evaluate models at “user language modeling” - how well they match the distribution of human language on the held-out test of conversations. To measure this, we compute model perplexity (PPL), a standard intrinsic metric commonly used to assess how well an LM predicts a sample of text Brown et al. (1992); Hewitt et al. (2023). For a test set that consists of a total of $N$ tokens, we compute the per-token perplexity for a model $P_{\theta}$ as $\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P_{\theta}(x_{i}\mid x_{<i})\right)$. Lower PPL reflects less surprisal and better alignment with the underlying data distribution. We evaluate on held-out test samples from WildChat and on the entire PRISM dataset Kirk et al. (2024), which we use as an out-of-domain test set and consists of 8,011 conversations. We applied the identical intent generation procedure to PRISM conversations.
@@ 3.1 Multi-Turn Interaction
Human users often omit details in their prompts, phrase requests in varied ways, and rely on the broader conversational context, expecting the assistant to interpret and leverage this information across turns Herlihy et al. (2024); Yang et al. (2025). It is therefore essential for our user LM to gradually progress through its intent across multiple turns, rather than revealing all the details in a single turn. We measure this by computing the average overlap of 1-grams between the generated user turns and the generic intent. We remove all stopwords before computing overlap. A lower overlap is particularly desirable because it indicates that the model expresses its intent using varied language while introducing details progressively.
@@
We assess the ability of each simulator to terminate the conversation when the conversation ended with the real user. To measure this, we compare the model’s generation of the <|endconversation|> token with actual conversation endings in PRISM conversations. Treating this as a binary classification task, we compute the F1 score to quantify how accurately the model predicts conversation endings. A high F1 score indicates that the model not only ends conversations at appropriate points but also avoids stopping too early or going on excessively.
@@ 3.2 Simulation Robustness
Text generated by a user LM should resemble the natural way humans write and be distinguishable from the typical style of assistant LMs. To measure the naturalness of the utterances, we use the state-of-the-art AI-detector Pangram Emi & Spero (2024), which was shown by Russell et al. (2025) to achieve near-perfect detection accuracy, on par with frequent users of ChatGPT. For our evaluation, we sample 2,000 first-turn generations, ensuring that each utterance ranges between 50 and 200 tokens, enough text for the detector to perform its prediction reliably. Pangram returns a likelihood for the text of being human-written, which we average across all 2,000 utterances.
@@
Assistants sometimes get confused in a conversation and, as an attempt to clarify, they suggest something the user did not want. Robust user simulators must adhere to their intent and not comply with assistant suggestions that steer away from their intent. To measure how well models adhere to their intent, we test them under the following setup. We initiate conversations where the user asks an open-ended question (e.g., who sang the song i’m a nut?). We then prompt GPT-4o to generate a response stating uncertainty of the answer and suggesting to assist with something else (e.g., I’m not sure what the answer is to who sang the song “I’m a Nut.” However, if you’re interested in exploring music, how about looking into the fascinating world of unusual instruments?). We then generate the next user turn for each model and measure the rate at which they refuse the assistant’s suggestion and adhere to their original intent. We perform this test using 2,000 random samples from the NaturalQuestions Kwiatkowski et al. (2019) as the first user turn. We evaluate whether the model adheres to its intent at the first turn by prompting GPT-4o as a judge.
@@ 4.1 Experimental Setting
We adopt the simulation setting of Laban et al. (2025), which simulates multi-turn conversations for various generation tasks. Specifically, we use 65 task intents that involve users completing a math word problem (based on GSM8k Cobbe et al. (2021)) or writing a Python program (based on HumanEval Chen et al. (2021)). Both tasks are appealing as they represent common use cases of current LLMs Tomlinson et al. (2025), and can be evaluated in a verifiable manner. We use UserLM-8b to simulate 10 conversations for each task intent (a total of 650 simulations). We also simulate conversations by prompting GPT-4o-mini and GPT-4o as the user simulators. In all conversations, we use GPT-4o as the assistant LM, keeping this aspect of the simulation fixed.
@@
We performed a quantitative analysis of the simulations, computing eight metrics focused on five aspects of the simulator: (1) Intent Coverage: measuring whether the simulator reveals the intent information adequately in the course of the conversation, (2) Information Diversity: evaluating how the user simulator repeats, skips, or adds information during conversations, (3) Pace Diversity: evaluating whether the user simulator can reveal information at varied speeds, effectively simulating users with differing communication styles, (4) Lexical Diversity: assessing the capacity of the simulator in using unique wording in each simulation, and (5) Assistant Performance: measuring the impact of the user simulator on the assistant’s ability to complete the task. The implementation details of each metric are provided in Appendix C.3.
@@ 4.2 Simulation Results & Analysis
The three simulators cover 76-86% of the units of information from the original intent, indicating they successfully stay on topic. Our UserLM-8b is more likely to repeat required information across turns, and less likely to reveal information that is not required to accomplish the task (e.g., revealing a trivial base case for a Python function). In contrast, the GPT-based user simulators are more monotonic: they tend to reveal information from the intent once, with less emphasis on revising what they mean by repetition or omitting information. UserLM-8b is the only simulator that introduces additional demands not specified in the original intent. We performed manual tagging of the additional demands specified by the UserLM-8b, with the three main categories being: (1) providing example test cases (34%, e.g., get_closest_vowel("FULL") should return U), (2) defining naming conventions (21%, e.g., The function should be named ‘rabbit’), and (3) implementation constraints (20%, e.g., Avoid using the built-in min and max functions). In contrast, the GPT-based simulators stick to the script, rarely injecting demands, leading to more homogeneous simulations.
@@
Our UserLM-8b achieves higher lexical diversity. In other words, two simulations based on the same intent have less lexical overlap when using UserLM-8b compared to using GPT-4o as the simulator. The diversity is due to UserLM-8b varying its language and style in different conversations, in contrast to GPT-based simulators that retain high similarity with the intent and avoid language variation.
@@ 5 Discussion & Implications
A key repeating insight from our results is that scaling up assistant LMs does not automatically result in improved user simulators. We first see this in Section 2.3 when comparing PPL on user utterances across instruction-tuned models, finding that Llama3-8b-Instruct exhibited worse PPL than the smaller 1B model. Similarly, in our analysis on initializing training from instruction-tuned checkpoints, starting from the 1b assistant model resulted in lower PPL (18.45) compared to starting from the 8b assistant model (27.25). This pattern also emerged in some of the intrinsic evaluations where GPT-4o-mini surpassed GPT-4o, particularly in terms of robustness, demonstrating a much better ability to adhere to its user role and intent. These results suggest that better assistant LMs do not make for better user simulators, highlighting the need to explicitly train user LMs. Looking at user LMs, our experiments show that scaling user LM training from 1b to 8b leads to gains across the board, with further scaling likely to be beneficial.
@@
In this work, user LMs were trained to simulate a broad, general audience, capturing behaviors commonly shared across a population of users. As such, UserLM-8b is a general-purpose foundation model that can model users in WildChat and PRISM. However, users with different demographics or personas may exhibit important variations that are important to model. For instance, non-native English speakers in Middle Eastern countries often phrase requests differently from native speakers in the United States, reflecting their own dialects and linguistic backgrounds. We believe a key focus for future studies on user language modeling is the development of more personalized user LMs that can simulate specific user groups or behavior in specific domains. The collection of large-scale data in specific settings is often challenging, and we hope the user LMs we propose can serve as a foundation to finetune more specialized user LMs, effectively lowering the data requirements in training personalized models.
@@
In Section 4, we presented one use case of user LMs in providing a more realistic evaluation of assistants in multi-turn conversations. Beyond interactive evaluation, we believe user LMs can open up a range of promising applications. One such direction is user modeling, where prior work has explored the use of LMs to estimate distributions of user responses to surveys Hu & Collier (2024); Suh et al. (2025). User LMs can be seen as a generative extension of this approach: instead of merely predicting distributions, they can generate natural text responses to survey questions, providing estimates of typical user behaviors. A second potential use case is in improving judge models. Current approaches often train judge models based on assistant LMs to approximate user preferences, but these are prone to assistant-specific biases and limited personalization Dong et al. (2024). By contrast, user LMs could serve as more realistic judges, either through prompting or finetuning into user LM judge models. Finally, user LMs can be leveraged for synthetic conversational data generation Wang et al. (2025). As we show, user LMs exhibit greater lexical diversity and produce more varied simulations, making them better-suited for generating synthetic data that can be used to fine-tune assistants and improve their robustness.
@@
The experiments we present in this work use automated methods to simulate users, presenting findings that are expected to generalize to real user behavior. However, simulating user behavior is a challenging problem, and we do not anticipate simulations to be applicable to all cases. In particular, working with experts in technical domains such as law Das et al. (2025), creative writing Chakrabarty et al. (2025) or science Asai et al. (2024) remains necessary. We propose the following scope: user simulation is a useful tool in scaling experiments, allowing the efficient discovery of system flaws that affect broad and generic populations, whereas working with experts or specific user populations is necessary to gain knowledge about user nuances.
More open-source base models will be needed.
Our experiments show that training user LMs from base LMs is more effective than starting from an instruction-tuned checkpoint, where the model has already been tuned to behave as an assistant. At the present time, most releases of open-weight LMs consist of instruction-tuned assistant models, with few releases of base models. We hope that our study encourages further research on building safe and general-purpose base models, as they will be useful to the development of user LMs.
@@
The user LMs we release as part of this work are relatively small (1b and 8b parameters) and have been post-trained on a limited set of natural user-assistant interactions (343,951 conversations). Despite this, they still outperform prompted assistants such as GPT-4o in simulating users and provide a more comprehensive evaluation of multi-turn assistant performance on downstream tasks. We expect that scaling up training and gaining access to higher-quality corpora of user logs will result in more performant user LMs, serving as more realistic simulation environments for assistants during reinforcement learning. This, in turn, will help develop more robust and effective assistants. We view our models as an initial milestone toward this long-term goal.

### v2

@@ Abstract
Conversations with LMs involve two participants: a human user leading the conversation, and an LM assistant responding to the user’s request. To satisfy this specific role, LMs are post-trained to be helpful assistants – optimized to produce exhaustive and well-structured responses, free of ambiguity and grammar errors. User utterances, on the other hand, are rarely perfected, with each user phrasing requests in unique ways, sometimes putting in partial effort at each turn and refining on the fly. To evaluate LM performance in realistic settings, prior work simulated users in multi-turn conversations, often by prompting an LM originally trained to be a helpful assistant to act as a user. However, we show that assistant LMs make for poor user simulators, with the surprising finding that better assistants yield worse simulators. Instead, we introduce purpose-built User Language Models (User LMs) - models post-trained to simulate human users in multi-turn conversations. Through various evaluations, we show how User LMs align better with human behavior and achieve better simulation robustness than existing simulation methods. When leveraging User LMs to simulate coding and math conversations, the performance of a strong assistant (GPT-4o) drops from 74.6% to 57.4%, confirming that more realistic simulation environments lead to assistant struggles as they fail to cope with the nuances of users in multi-turn setups.
@@ 2.1 Problem Description
Similar to how assistant LMs must follow instructions, user LMs must follow an intent that directs the conversation. Initial experimentation revealed that there exists a fine balance in defining user intent. On one extreme, not providing any intent at all reduces the usability of the simulator, as it cannot be effectively directed towards a task. On the other extreme, a fully-specified intent that contains all information of the user utterances renders the simulator obsolete: it simply parrots information from the intent with little added value. In this work, we define user intents as high-level conversation objectives: capturing the overall goal of the user without mentioning explicit details (see examples in Figure 2) – achieving a balance between the two extremes. Preliminary analyses showed that user LMs trained on high-level intents were practically more useful than ones trained without intent or with fully-specified intents: they can be guided to simulate specific conversations, but are in charge of language choices that allow for diverse and realistic simulation.
@@ 2.2 Experimental Details
We use WildChat Zhao et al. (2024), which contains 478,498 English conversations between ChatGPT and users from 192 countries. We noticed many near-duplicate samples and performed de-duplication, which resulted in 384,336 conversations (see details in Appendix A). The authors of WildChat did not define data splits (i.e., train vs. test); therefore, we created our own by first identifying unique users based on their hashed IP addresses and countries. For each country, we randomly split users into a 90/5/5 split for training, validation, and testing. All conversations from the same user were then assigned to the same split, ensuring that each split contains conversations from entirely distinct users. This resulted in 343,951 conversations for training, 22,442 for validation, and 17,943 for testing. After formatting conversations as we described in §2.1, we obtained 1,047,930 training samples, 118,291 validation samples, and 137,224 testing samples.
@@
We perform full-parameter fine-tuning of Llama3-8b-Base and Llama3.2-1b-Base. As we will show, we find that starting from the base model is better than an instruction-tuned checkpoint. We used a max sequence length of 2048 tokens, a batch size of 1024 samples, and a learning rate of 2e-5. Training was done on 4 A6000 GPUs, taking 62 hours to train the 1b model and 227 hours for the 8b model. We refer to our fine-tuned user LMs as UserLM-1b and UserLM-8b.
@@
We compare to multiple assistant LMs: Llama3.2-1b-Instruct, Llama3-8b-Instruct, and GPT-4o. Prompting assistant LMs to role-play users is the predominant approach in prior work on user simulation Chang et al. (2025); Ivey et al. (2024). We further engineered the prompt of the assistant-based simulators to give explicit instructions on how to replicate user behavior: splitting content across turns, not being verbose, or making occasional typos, without overdoing such phenomena. Starting from prompts used in prior work, we selected the best-performing variant based on validation perplexity on WildChat (see prompts and selection details in Appendix D). We also compare to USP-8b Wang et al. (2025), a fine-tuned version of Llama3-8b-Base on a subset of 94,874 LMSys-Chat conversations to generate user utterances Zheng et al. (2023). All baselines were instructed to respond with <|endconversation|> if they believe the user intent is fulfilled.
@@ 2.3 Initial Analyses
We evaluate models at “user language modeling” - how well they match the distribution of human language on the held-out test set of conversations. To measure this, we compute model perplexity (PPL), a standard intrinsic metric commonly used to assess how well an LM predicts a sample of text Brown et al. (1992); Hewitt et al. (2023). For a test set that consists of a total of $N$ tokens, we compute the per-token perplexity for a model $P_{\theta}$ as $\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P_{\theta}(x_{i}\mid x_{<i})\right)$. Lower PPL reflects less surprisal and better alignment with the underlying data distribution. We evaluate on held-out test samples from WildChat and on the entire PRISM dataset Kirk et al. (2024), which we use as an out-of-domain test set and consists of 8,011 conversations. We applied the identical intent generation procedure to PRISM conversations.
@@ 3.1 Multi-Turn Interaction
Human users often omit details in their prompts, phrase requests in varied ways, and rely on the broader conversational context, expecting the assistant to interpret and leverage this information across turns Herlihy et al. (2024); Yang et al. (2025). It is therefore essential for our user LM to gradually progress through its intent across multiple turns, rather than revealing all the details in a single turn. We measure this by computing the average overlap of 1-grams between the generated user turns and the generic intent for each conversation in PRISM. We remove all stopwords before computing overlap. Since the task intents use generic language that does not match the language of real user utterances (as confirmed by the very low n-gram overlap between test intents and real user utterances), a lower overlap indicates that the model rephrases intent information using its own varied language rather than copying verbatim from the intent, while introducing details progressively across turns.
@@
We assess the ability of each simulator to terminate the conversation when the conversation ended with the real user. To measure this, we compare the model’s generation of the <|endconversation|> token with actual conversation endings in the PRISM dataset. Treating this as a binary classification task, we compute the F1 score to quantify how accurately the model predicts conversation endings. A high F1 score indicates that the model not only ends conversations at appropriate points but also avoids stopping too early or going on excessively.
@@ 3.2 Simulation Robustness
Text generated by a user LM should resemble the natural way humans write and be distinguishable from the typical style of assistant LMs. To measure the naturalness of the utterances, we use the state-of-the-art AI-detector Pangram Emi & Spero (2024), which was shown by Russell et al. (2025) to achieve near-perfect detection accuracy, on par with frequent users of ChatGPT. For our evaluation, we sample 2,000 first-turn generations, ensuring each utterance ranges between 50 and 200 tokens, enough text for the detector to perform its prediction reliably. Pangram returns a likelihood of being human-written, which we average across all 2,000 utterances.
@@
Assistants sometimes get confused in a conversation and, as an attempt to clarify, they suggest something the user did not want. Robust user simulators must adhere to their intent and not comply with assistant suggestions that steer away from their intent. To measure this, we initiate conversations where the user asks an open-ended question (e.g., who sang the song i’m a nut?). We then prompt GPT-4o to generate a response stating uncertainty of the answer and suggesting to assist with something else (e.g., I’m not sure what the answer is to who sang the song “I’m a Nut.” However, if you’re interested in exploring music, how about looking into the fascinating world of unusual instruments?). We then generate the next user turn and measure the rate at which the simulator refuses the assistant’s suggestion and adhere to their original intent. We perform this test using 2,000 random samples from NaturalQuestions Kwiatkowski et al. (2019) as the first user turn. We evaluate whether the model adheres to its original intent by prompting GPT-4o as a judge.
@@ 4.1 Experimental Setting
We adopt the simulation setting of Laban et al. (2025), which simulates multi-turn conversations for various generation tasks. Specifically, we use 65 task intents that involve users completing a math word problem (based on GSM8k Cobbe et al. (2021)) or writing a Python program (based on HumanEval Chen et al. (2021)). Both tasks are appealing as they represent common use cases of current LLMs Tomlinson et al. (2025), and can be evaluated in a verifiable manner. We use UserLM-8b to simulate 10 conversations for each task intent (a total of 650 simulations). We also simulate conversations by prompting GPT-4o-mini and GPT-4o as the user simulators. In all conversations, we use GPT-4o as the assistant LM, keeping this aspect of the simulation fixed. For UserLM-8b, we applied a set of simple generation guardrails (detailed in Appendix C.1) to counteract noise from its smaller model size; these guardrails were only used in this section and do not affect the evaluations in Tables 1 and 2.
@@
We performed a quantitative analysis of the simulations, computing eight metrics focused on five aspects of the simulator: (1) Intent Coverage: whether the simulator reveals the intent information adequately in the course of the conversation, (2) Information Diversity: how the simulator repeats, skips, or adds information during conversations, (3) Pace Diversity: whether the simulator can reveal information at varied speeds, effectively simulating users with differing communication styles, (4) Lexical Diversity: the capacity of the simulator in using unique wording in each simulation, and (5) Assistant Performance: the impact of the simulator on the assistant’s ability to complete the task. The implementation details of each metric are provided in Appendix C.3.
@@ 4.2 Simulation Results & Analysis
The three simulators cover 76-86% of the units of information from the original intent, indicating they successfully stay on topic. Our UserLM-8b is more likely to repeat required information across turns, and less likely to reveal information that is not required to accomplish the task. In contrast, the GPT-based simulators are more monotonic: they tend to reveal information from the intent once, with less emphasis on revising what they mean by repetition or omitting information. UserLM-8b is the only simulator that introduces additional demands not specified in the original intent. We manually labeled the additional demands specified the UserLM-8b, with the three main categories being: (1) providing example test cases (34%, e.g., get_closest_vowel("FULL") should return U), (2) defining naming conventions (21%, e.g., The function should be named ‘rabbit’), and (3) implementation constraints (20%, e.g., Avoid using the built-in min and max functions). In contrast, the GPT-based simulators stick to the script, rarely injecting demands, leading to more homogeneous simulations.
@@
UserLM-8b achieves higher lexical diversity. In other words, two simulations based on the same intent have less lexical overlap when using UserLM-8b compared to using GPT-4o as the simulator. The diversity is due to UserLM-8b varying its language and style in different conversations, in contrast to GPT-based simulators that retain high similarity with the intent and avoid language variation.
@@ 5 Discussion & Implications
A repeating insight from our results is that scaling up assistant LMs does not automatically result in improved simulators. We first see this in §2.3 when comparing PPL on user utterances, finding that Llama3-8b-Instruct exhibited worse PPL than the smaller 1b model. Similarly, in our analysis on initializing training from instruction-tuned checkpoints, starting from the 1b assistant model resulted in lower PPL (18.45) compared to starting from the 8b assistant model (27.25). This pattern also emerged in some of the intrinsic evaluations where GPT-4o-mini surpassed GPT-4o, particularly in terms of intent and role adherence. These results highlight the need to explicitly train user LMs.
@@
In this work, user LMs were trained to simulate a broad, general audience, capturing behaviors commonly shared across a population of users. However, users with different demographics or personas may exhibit important variations that are important to model. For instance, non-native English speakers in Middle Eastern countries often phrase requests differently from native speakers in the United States, reflecting their own dialects and linguistic backgrounds. We believe a key focus for future studies is the development of more personalized simulators that can simulate specific user groups or behavior in specific domains. The collection of large-scale data in specific settings is often challenging, and we hope the user LMs we propose can serve as a foundation to finetune more personalized user LMs, lowering data requirements.
@@
Beyond interactive evaluation, we believe user LMs can open up a range of promising applications. One direction is user modeling, where prior work has explored the use of LMs to estimate distributions of user responses to surveys Hu & Collier (2024); Suh et al. (2025). User LMs can be seen as a generative extension of this approach: instead of predicting distributions, they can generate natural text responses to survey questions, providing estimates of user behaviors. A second use case is improving judge models. Current approaches often train judge models based on assistant LMs to approximate user preferences, but these are prone to assistant-specific biases and limited personalization Dong et al. (2024). By contrast, user LMs could serve as more realistic judges through prompting or finetuning into user LM judge models. Finally, user LMs can be leveraged for synthetic data generation Wang et al. (2025). As we show, user LMs produce more varied simulations, making them better-suited for generating synthetic data that can be used to fine-tune assistants and improve their robustness.
@@
The experiments we present in this work use automated methods to simulate users, presenting findings that are expected to generalize to real user behavior. However, simulating user behavior is a challenging problem, and we do not anticipate simulations to be applicable to all cases. In particular, working with experts in technical domains such as law Das et al. (2025), creative writing Chakrabarty et al. (2025) or science Asai et al. (2024) remains necessary. We propose the following scope: user simulation is a useful tool in scaling experiments, allowing the efficient discovery of system flaws that affect broad populations, whereas working with experts or specific user populations is necessary to gain knowledge about user nuances.
@@
The user LMs we release as part of this work are relatively small (1b and 8b parameters) and have been post-trained on a limited set of natural user-assistant interactions (343,951 conversations). Despite this, they still outperform prompted assistants such as GPT-4o in simulating users and provide a more comprehensive evaluation of multi-turn assistant performance on downstream tasks. We expect that scaling up model size and training data will result in better user LMs, serving as more realistic simulation environments. This, in turn, will help develop more robust and effective assistants.